OpenStudy (anonymous):

Can somebody help me find the value for 'X' in this equation? 10^x-7 = 10^3x+8

OpenStudy (mrnood):

is this the correct equation? \[10^{(x-7)}= 10^{(3x+8)}\] or \[10^{x}-7= 10^{3x}+8\]

OpenStudy (anonymous):

The first one sir.

OpenStudy (mrnood):

In that case - you can ignore the fact that they are exponents (x-7)=(3x+8)

OpenStudy (anonymous):

Okay, could you show me the method to solving for x?

OpenStudy (mrnood):

No - I won't show you that. If you are dealing with equations such as in your question then you must be able to solve the simple equation above. Simply move all x terms to one side, and constant terms to the other - taking note of the signs....

OpenStudy (anonymous):

I'm sorry i'm still having a little trouble putting in the equation. Im not the best at this. Could you show me how the equation is the put together to solve?

OpenStudy (mrnood):

I'm sorry - I believe that you must have the skills to solve this yourself x-7 = 3x+8

OpenStudy (mrnood):

Add 7 to each side of the equation Then subtract 3x from each side of the equation

OpenStudy (anonymous):

After that would the new equation would be x-11 = 7x+8 ?

OpenStudy (mrnood):

start by adding 7 to each side of the equation

OpenStudy (anonymous):

X-14 = 10x+8?

OpenStudy (mrnood):

write the equation out as it is - but add "+7" to each side

OpenStudy (anonymous):

x-7+7=3x+8+7

OpenStudy (mrnood):

OK - now do the adding up of JUST the numbers on each side

OpenStudy (anonymous):

x-14=3x+15?

OpenStudy (mrnood):

Not quite right The right side is correct - but the left side is wrong x-7+7=3x+8+7 Try again

OpenStudy (anonymous):

x7=3x+15

OpenStudy (anonymous):

the 7 becomes positive?

OpenStudy (mrnood):

sorry - you started by writing an equation involving exponents, but now cannot add -7+7 Are you focussed on doing this?? - I will not give oyu the answer...

OpenStudy (anonymous):

Yeah, I am.

OpenStudy (mrnood):

so - what is x-7+7 ??

OpenStudy (anonymous):

0?

OpenStudy (mrnood):

no start with x ... subtract 7 .... add 7 ....... what is left?

OpenStudy (anonymous):

Im getting zero.

OpenStudy (mrnood):

Have another think - when you don't get 0 come back

OpenStudy (mrnood):

hint: x represents some unknown number - any unknown number.....

OpenStudy (anonymous):

3?

OpenStudy (mrnood):

I don't think oyu are in a position to proceed with this question. I can't see how oyu are asking questions of this sort without having already been able to solve an equation such as you have. I don't mean to discourage you - but you must work on some basic algebra (equations) before you can tackle this. Good luck

OpenStudy (anonymous):

Okay. Thanks for the help anyways.